# When Does Divide and Conquer Work for Long Context LLM? A Noise Decomposition Framework

### v1

@@ Abstract
We investigate the challenge of applying Large Language Models (LLMs) to long texts. We propose a theoretical framework that distinguishes the failure modes of long context tasks into three categories: cross-chunk dependence (task noise), confusion that grows with context size (model noise), and the imperfect integration of partial results (aggregator noise). Under this view, we analyze when it is effective to use multi-agent chunking, i.e., dividing a length sequence into smaller chunks and aggregating the processed results of each chunk. Our experiments on tasks such as retrieval, question answering, and summarization confirm both the theoretical analysis and the conditions that favor multi-agent chunking. By exploring superlinear model noise growth with input length, we also explain why, for large inputs, a weaker model configured with chunk-based processing can surpass a more advanced model like GPT4o applied in a single shot. Overall, we present a principled understanding framework and our results highlight a direct pathway to handling long contexts in LLMs with carefully managed chunking and aggregator strategies [^1] The code implementation will be available at https://github.com/NehzUx/DivideConquerAgents.
@@ 1 Introduction
In this paper, we propose a divide and conquer framework Qian et al. (2024b); Zhang et al. (2024b); Zhou et al. (2024) that offers a more systematic view at these issues. we analyze these challenges by establishing a theoretical framework that tracks three main sources of error in long-context tasks. First, task noise arises from cross-chunk dependencies that cannot be handled by processing each segment alone. Second, model noise arises from degraded performance with increasing input length. Third, aggregator noise emerges when partial results are combined incorrectly, even if each chunk was handled well. We show how the balance of these factors determines when chunk-based division is advantageous, and how the design of the aggregator stage plays an important role. In particular, when a task demands only mild cross-chunk reasoning, chunking can reduce model confusion with little downside. However, when cross-chunk synergy is large, only a more advanced aggregator can preserve important global connections.
We also show that model noise can worsen at a rate that outstrips the benefit of seeing the entire input at once. Beyond a certain length threshold, splitting the text can yield better performance, even if the individual worker models are weaker than a single large model. We attribute this to a superlinear increase in model error growth with input length. Our experiments on retrieval, QA, summarization, and other tasks confirm that this superlinear effect surfaces in practice, and demonstrate how a planner can arrange prompts to keep aggregator noise manageable. Across these long context tasks, weak models handling chunks often outperform a single-shot strong model (e.g. GPT4o).
@@
  • We provide empirical evidence of superlinear performance degradation when input length is large. We then demonstrate that splitting into smaller pieces can help mitigate these effects, except in the case where cross-chunk synergy is very high.
@@ 2.1 Long context LLM
Developing long context LLMs has seen increasing popularity in applications such as long document understanding. The transformer architecture plays a dominating role in language modeling and, due to the attention mechanism, it is associated with quadratic computational and memory complexity against sequence length(Tay et al., 2022). Therefore, it is expensive to train and serve long context LLMs in practice. To mitigate this, researchers have explored various efficient transformer architectures, such as block wise attention(Qiu et al., 2020), window attention(Beltagy et al., 2020), low-rank approximated attention(Wang et al., 2020), routing attention(Kitaev et al., 2020), etc. (Dao et al., 2022) proposed flash attention and (Liu et al., 2023) designed ring attention, both contributed to efficient training of long context LLMs. The recent progress on transformer alternatives also show promises of overcoming the computational complexity of attention mechanisms, including state-space models, e.g, Mamba(Gu and Dao, 2024) and convolutional-based approach, Hyena(Poli et al., 2023). On the other hand, given existing LLMs trained on short context, many research works have been presented on length extrapolation based on positional encoding(Chen et al., 2023; Peng et al., 2024; Jin et al., 2024), and retrieval-augmented generation(Lewis et al., 2020; Wang et al., 2024b), where additional design complexity is be added on positional encoding adaptation and retriever.
@@ 2.2 Multi-Agents
Recent research has explored the potential of multi-agent systems utilizing large language models (LLMs) for complex task-solving. In the realm of task decomposition and planning, Qian et al. (2023) proposed a framework where multiple LLM-based agents collaboratively break down tasks into subtasks and develop execution plans. Similarly, Haji et al. (2024) introduced a hierarchical mechanism where tree of thoughts reasoners are combined with thought verifiers. Zhao et al. (2024b) breaks down and process long texts using multiple agents, enabling efficient handling of 128K token contexts through task decomposition.
Beyond task breakdown and planning, recent work focused on communication mechanisms within multiple agents. Several studies (Guo et al., 2024) have investigated various approaches to enable effective information exchange between agents. Common communication paradigms such as debate is commonly used (Du et al., 2024; Chan et al., 2024; Liang et al., 2023). Alternatively, Qian et al. (2024a) adapts a cooperative paradigm where agents work together towards a common goal. Many multi-agent works use a layered structure. Wang et al. (2024a) developed a multi-layered framework that separates LLMs into proposers and aggregators. Liang et al. (2023) encourages more divergent thinking though layered debate among debtors. However, a centralized sharing structure has also been explored by Hong et al. (2023).
@@ 2.3 Divide and conquer
Divide-and-conquer is a classic computational strategy that has recently been integrated with LLM-based multi-agent systems to handle long-context processing. Recent works such as LC-Boost Qian et al. (2024b), Chain-of-Agents (CoA) Zhang et al. (2024b), LongAgent Zhao et al. (2024a), and LLM×$\times$×MapReduce Zhou et al. (2024) , adopt this paradigm by splitting long inputs into smaller chunks processed by worker agents and then merging their outputs through a manager agent. These methods improve efficiency and enable long-text reasoning without requiring large context models.
However, existing approaches lack a formal theoretical framework to analyze the interaction between task complexity, model noise, and aggregation errors, making it difficult to optimize chunking strategies. They also struggle with understanding how cross-chunk dependencies impact performance, often leading to loss of contextual coherence when aggregating local outputs. Additionally, agent coordination mechanisms in prior work are heuristic-driven, lacking a principled way to balance worker and manager responsibilities. This work addresses these gaps by providing a rigorous modeling framework for analyzing noise propagation in Divide-and-conquer architectures, offering a systematic approach to optimizing multi-agent collaboration for long-context tasks.
@@ 3 Theoretical modeling framework
3.1 Problem statement and setting
We consider a long input $x$ of length $T$ tokens, denoted $x\in\mathcal{X}$, where $\mathcal{X}$ is the input space. The input $x$ can be split into $n$ contiguous chunks: $x\;=\;(x_{1},x_{2},\dots,x_{n}),\quad\text{with}\quad\sum_{i=1}^{n}\|x_{i}\|\;=\;T.$ A task asks for a mapping function $f:\mathcal{X}\to\mathcal{Y}$ where $\mathcal{Y}$ is the answer space. The target task function $y=f^{*}(x)$ is the correct or desired output for the entire input, for example returning the correct answer to a user query or extracting complex reasoning from a text of length $T$.
Single-shot long context agent inference.
We feed the entire $x$ (length $T$) into one agent that has context length equal to or longer than $T$. It computes $f_{\mathrm{single}}(x)\;=\;f^{*}(x)\;+\;\phi_{\mathrm{model}}(x),$ where $\phi_{\mathrm{model}}(x):\mathcal{X}\to\mathcal{Y}$ is a small-phi noise term capturing model errors (for example attention constraints or distribution mismatch). We denote $\phi_{\mathrm{model}}(x)=\phi(\|x\|)$, with $\phi(\cdot)$ non-decreasing. In practice, performance can degrade with $T$ even if $T$ is within the nominal context window.
Divide and conquer agents (D&C).
We split $x$ into $n$ subproblems $(x_{1},\dots,x_{n})$. Each subproblem is handled by a worker agent model $f_{i}(\cdot)$: $f_{i}(x_{i})\;=\;f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i}).$ For each chunk $x_{i}$, the worker agent noise is $\phi_{\mathrm{model}}(x_{i})$. If $\|x_{i}\|=\tfrac{T}{n}$ and $\phi(\cdot)$ grows faster than linearly (shown later), then single-shot usage over $T$ can be significantly noisier than chunk usage over $\tfrac{T}{n}$. Next, an aggregator (manager agent) $h(\cdot)$ merges these sub-results:
$f_{\mathrm{D\&C}}(x;n)$ $=h\bigl{(}f_{1}(x_{1}),\dots,f_{n}(x_{n})\bigr{)}=\prod_{i=1}^{n}f_{i}(x_{i})+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n}).$
If the aggregator $h$ is not fully capable of reconstructing global interactions, an additional noise $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ appears in the final output.
3.2 Task noise, model noise, and aggregator noise
We posit that many tasks can be partially factorized across chunks plus a noise term. Let
$g(x_{1},\dots,x_{n})\;=\;\underbrace{\prod_{i=1}^{n}g^{*}(x_{i})}_{\text{factorized component}}\;+\;\underbrace{\phi_{\mathrm{task}}(x_{1},\dots,x_{n})}_{\begin{subarray}{c}\text{cross-chunk dependency}\\ \text{or synergy}\end{subarray}},$
where $g^{*}(x_{i})$ is the local contribution from chunk $x_{i}$, and $\phi_{\mathrm{task}}(\cdot)$ measures how much $g$ departs from being a product of local chunk functions. If $\phi_{\mathrm{task}}\approx 0$, the task is nearly chunk-independent. Otherwise, large $\phi_{\mathrm{task}}$ indicates significant cross-chunk synergy.
When $\|\phi_{\mathrm{task}}\|\leq\epsilon$, a well-aligned aggregator $h$ can keep $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ on the order of $\epsilon$. Conversely, if $\phi_{\mathrm{task}}$ is large and the aggregator does not reprocess the entire $x$, then $\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$ may scale with $\|\phi_{\mathrm{task}}\|$.
Putting it all together. Summarizing the contributions of each term in D&C:
$f_{\mathrm{D\&C}}(x;n)$ $=\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})$
$=\prod_{i=1}^{n}f_{i}^{*}(x_{i})+\phi_{\mathrm{agg}}(x_{1},\dots,x_{n})+\Bigl{[}\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i})\Bigl{]}$
For simplicity, we then define $\Phi_{\mathrm{model}}:=\Bigl{[}\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i})\Bigl{]},\Phi_{\mathrm{agg}}:=\phi_{\mathrm{agg}}(x_{1},\dots,x_{n}),\Phi_{\mathrm{task}}:=-\phi_{\mathrm{task}}(x_{1},\dots,x_{n}).$ $\Phi_{\mathrm{model}}$ arises from model noise (and is a polynomial in the terms $\phi_{\mathrm{model}}(x_{i})$). $\Phi_{\mathrm{agg}}$ arises from aggregator limitations. $\Phi_{\mathrm{task}}$ is the negative of the task factorization mismatch.
Hence, we have $f_{\mathrm{D\&C}}(x;n)=\prod_{i=1}^{n}f_{i}^{*}(x_{i})+\Phi_{\mathrm{model}}+\Phi_{\mathrm{agg}}.$ We can compare it to the target $g(x)=\prod_{i=1}^{n}g^{*}(x_{i})+\phi_{\mathrm{task}}(x_{1},\dots,x_{n}).$ Then the approximation error is
$f_{\mathrm{D\&C}}(x;n)-g(x)$ $=\underbrace{\Bigl{[}\prod_{i=1}^{n}f_{i}^{*}(x_{i})-\prod_{i=1}^{n}g^{*}(x_{i})\Bigr{]}}_{\Delta^{*}(x)}\quad+\bigl{(}\,\Phi_{\mathrm{model}}\;+\;\Phi_{\mathrm{agg}}\;+\;\Phi_{\mathrm{task}}\,\bigr{)}.$ (1)
$\Delta^{*}(x)$ is the difference between the product of ideal worker outputs and the product of true chunk-level targets. If each $f_{i}^{*}(x_{i})$ matches $g^{*}(x_{i})$ well, the main discrepancy is $\Phi_{\mathrm{model}}+\Phi_{\mathrm{agg}}+\Phi_{\mathrm{task}}$. By improving worker models (to shrink $\phi_{\mathrm{model}}$) and using better aggregation (to reduce $\phi_{\mathrm{agg}}$), D&C can lower overall error. However, a large $\phi_{\mathrm{task}}$ still requires effective cross-chunk merging.
3.3 Why weaker models can outperform stronger ones
We first make precise the statement that $\phi_{\mathrm{model}}(x)$ grows faster than linearly in the input length, eliminating linear or sublinear growth cases:
Assumption 3.1 (Superlinear model noise).
Denote $\phi_{\mathrm{model}}(x)$ the model noise for an input of length $L=\|x\|$. For sufficiently large $L$, there exist constants $c>0$ and $\alpha>1$ such that $\phi_{\mathrm{model}}(x)\geq cL^{\alpha}.$
Justification by elimination. Please refer to Appendix A for details.
Theorem 3.2 (When D&C agents win).
Consider an input $x$ of length $T$. Suppose that the single-shot noise $\phi_{\mathrm{model}}(T)$ grows superlinearly, and that $\Delta^{*}(x)$, the aggregator noise $\Phi_{\mathrm{agg}}$, and the cross-chunk mismatch $\Phi_{\mathrm{task}}$ all remain bounded when the input is divided into $n$ chunks of length at most $L_{0}$. Then, for sufficiently large $T$, one can choose $n\approx T/L_{0}$ to ensure: $|f_{\mathrm{D\&C}}(x;n)-g(x)|<\phi_{\mathrm{model}}(T).$
Proof. Please refer to Appendix B for details.
3.4 Three regimes of noises
In many practical pipelines, $\Phi_{\mathrm{agg}}$ can be kept modest through careful prompt design by a planner agent. Thus, the major interplay often arises between $\Phi_{\mathrm{model}}$ (how confused the model is on each chunk) and $\Phi_{\mathrm{task}}$ (how strongly the chunks depend on each other). We characterize three regimes.
  • Case 1: Both $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$ are negligible. Here, each chunk is easy for the model, and the task factorization mismatch is minor. Splitting into chunks or using a single shot yields similar performance. Even if there is a slight aggregator noise $\Phi_{\mathrm{agg}}$, it stays negligible compared to the small $\Phi_{\mathrm{model}}$ and $\Phi_{\mathrm{task}}$.
  • Case 2: $\Phi_{\mathrm{task}}$ dominates. The input’s synergy is extensive, yet each individual chunk is straightforward. A naive aggregator might fail to merge the global context effectively, causing $\Phi_{\mathrm{agg}}$ to blow up unless prompted carefully. In this regime, single-shot usage could still do well if the model can handle long context, but D&C requires a robust aggregator for small $\Phi_{\mathrm{agg}}$.
  • Case 3: $\Phi_{\mathrm{model}}$ dominates. The task is nearly chunk-independent, but each chunk is intrinsically confusing to the model when taken in a single large context. Splitting the input reduces the per-chunk confusion, lowering $\Phi_{\mathrm{model}}$ significantly. Even if $\Phi_{\mathrm{agg}}$ grows somewhat, it is outweighed by the model noise improvement. This is precisely the domain where Thm 3.2 shows that D&C can outperform a single-shot approach.
These three cases illustrate that neither $\Phi_{\mathrm{task}}$ nor $\Phi_{\mathrm{model}}$ alone determines whether multi-agent chunking or single-shot usage is more suitable. Their relative magnitudes, together with $\Delta^{*}(x)$ and the aggregator’s effectiveness ($\Phi_{\mathrm{agg}}$), ultimately decide which approach is advantageous.

### v2

@@ Abstract
We investigate the challenge of applying Large Language Models (LLMs) to long texts. We propose a theoretical framework that distinguishes the failure modes of long context tasks into three categories: cross-chunk dependence (task noise), confusion that grows with context size (model noise), and the imperfect integration of partial results (aggregator noise). Under this view, we analyze when it is effective to use multi-agent chunking, i.e., dividing a lengthy sequence into smaller chunks and aggregating the processed results of each chunk. Our experiments on tasks such as retrieval, question answering, and summarization confirm both the theoretical analysis and the conditions that favor multi-agent chunking. By exploring the accelerated decay of model fidelity with input length, we also explain why, for large inputs, a weaker model configured with chunk-based processing can surpass a more advanced model like GPT4o applied in a single shot. Overall, we present a principled understanding framework and our results highlight a direct pathway to handling long contexts in LLMs with carefully managed chunking and aggregator strategies.
@@ 1 Introduction
In this paper, we propose a divide and conquer framework Qian et al. (2024b); Zhang et al. (2024c); Zhou et al. (2024) that offers a more systematic view at these issues. We analyze these challenges by establishing a theoretical framework that tracks three main sources of error in long-context tasks. First, task noise arises from cross-chunk dependencies that cannot be handled by processing each segment alone. Second, model noise arises from degraded performance with increasing input length. Third, aggregator noise emerges when partial results are combined incorrectly, even if each chunk was handled well. We show how the balance of these factors determines when chunk-based division is advantageous, and how the design of the aggregator stage plays an important role. In particular, when a task demands only mild cross-chunk reasoning, chunking can reduce model confusion with little downside. However, when cross-chunk synergy is large, only a more advanced aggregator can preserve important global connections.
We also show that model noise can worsen at a rate that outstrips the benefit of seeing the entire input at once. Beyond a certain length threshold, splitting the text can yield better performance, even if the individual worker models are weaker than a single large model. We attribute this to the superlinear growth of the model’s underlying length-induced degradation with input length. Our experiments on retrieval, QA, summarization, and other tasks confirm that this superlinear effect surfaces in practice, and demonstrate how a planner can arrange prompts to keep aggregator noise manageable. Across these long context tasks, weak models handling chunks often outperform a single-shot strong model.
@@
  • We provide empirical evidence of accelerated performance degradation when input length is large. We then demonstrate that splitting into smaller pieces can help mitigate these effects, except in the case where cross-chunk synergy is very high.
@@ 2.1 Long context LLM
Developing long context LLMs has seen increasing popularity in applications such as long document understanding. Due to the attention mechanism, transformer architecture is associated with quadratic computational and memory complexity against sequence length(Tay et al., 2022), thus expensive to train and serve in practice. To mitigate this, researchers have explored various efficient transformer architectures, such as block wise attention(Qiu et al., 2020), window attention(Beltagy et al., 2020), routing attention(Kitaev et al., 2020), etc. (Dao et al., 2022) proposed flash attention and (Liu et al., 2023) designed ring attention, both contributed to efficient training of long context LLMs. The recent progress on transformer alternatives also show promises of overcoming the computational complexity of attention mechanisms, including Mamba(Gu and Dao, 2024) and Hyena(Poli et al., 2023). Also, given existing LLMs trained on short context, many research works have been presented on length extrapolation based on positional encoding(Chen et al., 2023; Peng et al., 2024; Jin et al., 2024), and retrieval-augmented generation(Lewis et al., 2020; Wang et al., 2024b), where additional design complexity is added on positional encoding adaptation and retriever. Besides, a complementary line of work aligns LLMs for long contexts via preference/reward optimization, including LOGO (Tang et al., 2024), LongReward (Zhang et al., 2024a), and LongPO (Chen et al., 2025).
@@ 2.2 Multi-Agents
Recent research has explored the potential of multi-agent systems utilizing large language models (LLMs) for complex task-solving. In the realm of task decomposition and planning, Qian et al. (2023) proposed a framework where multiple LLM-based agents collaboratively break down tasks into subtasks and develop execution plans. Similarly, Haji et al. (2024) introduced a hierarchical mechanism where tree of thoughts reasoners are combined with thought verifiers. Zhao et al. (2024a) breaks down and process long texts using multiple agents, enabling efficient handling of 128K token contexts through task decomposition.
Beyond task breakdown and planning, recent work focused on communication mechanisms within multiple agents. Several studies (Guo et al., 2024) have investigated various approaches to enable effective information exchange between agents. Common communication paradigms such as debate is commonly used (Du et al., 2024; Chan et al., 2024; Liang et al., 2023). Alternatively, Qian et al. (2024a) adapts a cooperative paradigm where agents work together towards a common goal. Many multi-agent works use a layered structure. Wang et al. (2024a) developed a multi-layered framework that separates LLMs into proposers and aggregators. Liang et al. (2023) encourages more divergent thinking though layered debate among debaters. However, a centralized sharing structure has also been explored by Hong et al. (2023).
@@ 2.3 Divide and conquer
[…]
However, existing approaches lack a formal theoretical framework to analyze the interaction between task complexity, model noise, and aggregation errors, making it difficult to optimize chunking strategies. They also struggle with understanding how cross-chunk dependencies impact performance, often leading to loss of contextual coherence when aggregating local outputs. This work addresses these gaps by providing a rigorous modeling framework for analyzing noise propagation in Divide-and-conquer architectures, offering a systematic approach to optimizing multi-agent collaboration for long-context tasks.
@@ 3 Theoretical Modeling Framework
3.1 Problem Formulation: The Fidelity Decomposition
We model the divide-and-conquer (D&C) pipeline as an information transmission channel. Let $x$ be a long input of length $T$ split into $n$ contiguous chunks. Let $S(\hat{y})\in(0,1]$ represent the normalized evaluation metric for a prediction $\hat{y}$. We assume $S(\cdot)>0$ to ensure well-defined logarithmic losses.
We define the system’s score fidelity $\rho\in(0,1]$ as the ratio of the retained score to the ideal score. To analyze the error accumulation, we utilize the following telescoping identity, which decomposes the system fidelity into a product of score ratios in three stages (explained in following sections):
$\rho_{\mathrm{sys}}=\frac{S(h(\mathbf{\hat{a}}))}{S(y^{*})}=\underbrace{\frac{S(h^{*}(\mathbf{a}^{*}))}{S(y^{*})}}_{\rho_{\mathrm{task}}}\times\underbrace{\frac{S(h(\mathbf{a}^{*}))}{S(h^{*}(\mathbf{a}^{*}))}}_{\rho_{\mathrm{agg}}}\times\underbrace{\frac{S(h(\mathbf{\hat{a}}))}{S(h(\mathbf{a}^{*}))}}_{\rho_{\mathrm{model}}}.$ (1)
For analytical convenience, we analyze the system in log-space by defining the fidelity loss $\mathcal{L}\coloneqq-\log(\rho)$. The multiplicative nature of fidelity implies additive losses:
$\mathcal{L}_{\mathrm{sys}}=\mathcal{L}_{\mathrm{task}}+\mathcal{L}_{\mathrm{agg}}+\mathcal{L}_{\mathrm{model}}.$ (2)
We also define the total system error as $\mathcal{E}_{\mathrm{total}}\coloneqq 1-\rho_{\mathrm{sys}}$. In the high-fidelity regime, a first-order expansion yields $\mathcal{E}_{\mathrm{total}}\approx(1-\rho_{\mathrm{task}})+(1-\rho_{\mathrm{agg}})+(1-\rho_{\mathrm{model}})$ (Appendix B).
3.2 Stage 1: Decomposition (Task Fidelity)
Let $y^{*}=f^{*}(x)$ be the global ground truth. Let $\mathbf{a}^{*}=(a_{1}^{*},\dots,a_{n}^{*})$ denote the optimal chunk-level artifacts constrained by a fixed schema (e.g., restricted bandwidth or token budget). Specifically, $\mathbf{a}^{*}$ represents the information bottleneck imposed strictly by the decomposition interface. We define an ideal aggregator $h^{*}$ as the optimal function in the model class $\mathcal{H}$ given these artifacts:
$h^{*}=\arg\max_{h\in\mathcal{H}}S(h(\mathbf{a}^{*})).$
Task fidelity measures the information retained under these decomposition constraints:
$\rho_{\mathrm{task}}=\frac{S(h^{*}(\mathbf{a}^{*}))}{S(y^{*})}.$ (3)
If $\rho_{\mathrm{task}}\ll 1$, the task inherently resists decomposition under the chosen schema.
3.3 Stage 2: Aggregation (Aggregator Fidelity)
Let $h$ be the actual aggregator model used in the pipeline. We define Aggregator Fidelity by comparing the actual aggregator against the ideal aggregator on the same perfect inputs $\mathbf{a}^{*}$:
$\rho_{\mathrm{agg}}=\frac{S(h(\mathbf{a}^{*}))}{S(h^{*}(\mathbf{a}^{*}))}.$ (4)
This term isolates the limitation of the aggregator model itself (e.g., failure to reason over long sequences of artifacts), assuming perfect input artifacts.
3.4 Stage 3: Local Processing (Model Fidelity)
Let $\mathbf{\hat{a}}$ denote the actual noisy outputs generated by worker models. We define Model Fidelity as the ratio of performance given actual noisy inputs versus perfect inputs:
$\rho_{\mathrm{model}}=\frac{S(h(\mathbf{\hat{a}}))}{S(h(\mathbf{a}^{*}))}.$ (5)
We assume monotonicity in expectation: replacing noisy worker outputs with ideal artifacts does not degrade performance, ensuring $\rho_{\mathrm{model}}\leq 1$. The term $\mathcal{L}_{\mathrm{model}}$ quantifies the loss caused strictly by worker errors.
3.5 The D&C Advantage
We now formally characterize when a system of weaker agents outperforms a single stronger agent on very long inputs.
Proposition 3.1 (The D&C Advantage).
Let $\mathcal{L}_{\mathrm{strong}}(T)$ be the loss of a single strong model and $\mathcal{L}_{\mathrm{D\&C}}(T)$ be the loss of a divide-and-conquer system on input length $T$. Assume:
  1. 1.
    Super-Linear Collapse: The strong model’s loss grows super-linearly with context length: $\mathcal{L}_{\mathrm{strong}}(T)=\omega(T)$ (i.e., $\lim_{T\to\infty}\mathcal{L}_{\mathrm{strong}}(T)/T=\infty$).
  2. 2.
    Bounded Unit Loss: The D&C system processes inputs in fixed-size chunks, and the error per chunk (and associated overhead) is bounded by a constant.
Then, the D&C loss accumulates linearly ($\mathcal{L}_{\mathrm{D\&C}}(T)=O(T)$), and there exists a critical threshold $T_{0}$ such that for all $T>T_{0}$, the D&C system strictly outperforms the single strong model.
This proposition justifies D&C strategies despite the overhead: as $T\to\infty$, the super-linear "brain fog" of a single model inevitably exceeds the linear costs of decomposition and aggregation.
3.6 Three Regimes of Error
Based on the decomposition, we categorize long-context tasks into three regimes:
  • Regime 1: Trivial (Negligible Noise). $\mathcal{L}\approx 0$. Tasks like sparse retrieval where decomposition is lossless and models are robust.
  • Regime 2: The “Silo Effect” (Task Noise Dominates). $\mathcal{L}_{\mathrm{task}}\gg\mathcal{L}_{\mathrm{model}}$. The task requires global reasoning that is lost under the chunking schema. In this regime, D&C strategies saturate below the optimal performance regardless of model quality.
  • Regime 3: The “Brain Fog” (Model Noise Dominates). $\mathcal{L}_{\mathrm{model}}\gg\mathcal{L}_{\mathrm{task}}$. The input length $T$ is so massive that single-agent fidelity collapses. This is the optimal regime for D&C strategies.